OpenStudy (anonymous):

CeeCee recorded the number of necklaces for sale at 10 different locations and then arranged the numbers from smallest to largest. The numbers of necklaces for sale at the locations were 4, 4, 4, 5, 8, 9, 9, 9, 9, and 9.The mean number of necklaces for sale at the locations was 7. What is the mean absolute deviation? A. 0 B. 2 C. 2.2 D. 4.4

OpenStudy (anonymous):

any ideas?

OpenStudy (anonymous):

add all the numbers it gives you and divide by 10 then subtract whatever you get here ^ with each number. add all these numbers ^ divide by 10 answer^

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

will u check me

OpenStudy (anonymous):

7 is the average

OpenStudy (anonymous):

right

OpenStudy (anonymous):

4+7=11 4+7=11 4+7=11 5+11=16 8+11=19 9+11=20 9+11=20 9+11=20 9+11=20 9+11=20 Now what?

OpenStudy (anonymous):

wait, can you explain why you did that?

OpenStudy (anonymous):

i got 7as my average and added all of the numbers by it

OpenStudy (anonymous):

4+7=11 4+7=11 4+7=11 5+7= 8+7= 9+7= 9+7= 9+7= 9+7= 9+7=

OpenStudy (anonymous):

i did that

OpenStudy (anonymous):

add 7 to them all, not 11

OpenStudy (anonymous):

sorry i got 11 as 7+4 and got confused

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

5+7=2 8+7=15 9+7=16 16 16 16 16

OpenStudy (anonymous):

ok what next

OpenStudy (anonymous):

add all those numbers together, then divide by 10

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

done

OpenStudy (anonymous):

not yet im adding

OpenStudy (anonymous):

is 12.8 correct?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

its wrong

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

idk then

OpenStudy (anonymous):

can u check it

OpenStudy (anonymous):

please